OpenStudy (anonymous):

perform the indicated operation and simplify 2/x-4 + 1/x+6

jimthompson5910 (jim_thompson5910):

\[\large \frac{2}{x-4} + \frac{1}{x+6}\] \[\large \frac{2(x+6)}{(x-4)(x+6)} + \frac{1}{x+6}\] \[\large \frac{2(x+6)}{(x-4)(x+6)} + \frac{1(x-4)}{(x-4)(x+6)}\] \[\large \frac{2x+12}{(x-4)(x+6)} + \frac{x-4}{(x-4)(x+6)}\] \[\large \frac{2x+12+x-4}{(x-4)(x+6)}\] \[\large \frac{3x+8}{(x-4)(x+6)}\] \[\large \frac{3x+8}{x^2+2x-24}\] So \[\large \frac{2}{x-4} + \frac{1}{x+6}=\frac{3x+8}{x^2+2x-24}\]